OpenStudy (anonymous):

Rocky simplified an expression in three steps, as shown x to the power of negative 5 multiplied by y to the power of 2, over y multiplied by x to the power of 3 multiplied by x to the power of 3 multiplied by y to the power of negative 5, the whole to the power of 2 equals x to the power of negative 10 multiplied by y to the power of 4, over y to the power of 2 multiplied by x to the power of 6 multiplied by x to the power of 6 multiplied by y to the power of negative 10 equals x to the power of negative 10 multiplied by y to the power of 4, over y to the power of negative 8 multiplied by x

OpenStudy (anonymous):

PLEASE HELP ME I NEED TO PASS

OpenStudy (anonymous):

@mathstudent55 @robtobey @freckles @mathmate @texaschic101

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

hello??

OpenStudy (anonymous):

thats what i chose

OpenStudy (anonymous):

The final answer is:\[\frac{y^{12}}{x^{22}} \]

OpenStudy (anonymous):

In Step 2, y^-8 is 1/y^8 . When inverted and multiplied by the numerator the exponent of y in the numerator becomes 12. The same idea applies to x^-10 in the numerator. Can you see why on you own analysis?

OpenStudy (anonymous):

Refer to the attachment from the Mathematica 9 program.

OpenStudy (anonymous):

ok i need help with one more

OpenStudy (anonymous):

Which of the following is a step in simplifying the expression x multiplied by y to the power of 3 over x to the power of negative 4 multiplied by y to the power of 4, the whole to the power of negative 2.? x to the power of negative 2 multiplied by y to the power of negative 6, the whole over x to the power of 8 multiplied by y to the power of negative 8. x to the power of negative 2 multiplied by y, the whole over x to the power of negative 6 multiplied by y to the power of 2. x to the power of negative 2 multiplied by y to the power of negative 6, the whole over x to the power of negative 4 multiplied by y to the power of 4. x to the power of negative 2 multiplied by y, the whole over x to the power of negative 4 multiplied by y to the power of 4.

OpenStudy (anonymous):

Hint:\[\frac{1}{\left(\frac{a}{b}\right)^2}=\frac{b^2}{a^2}\]

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

what do you think the answer is i need to pass 8th grade and this is my final test